(*) If A,B are acute angles, tan A = n/(n+1), tan B = 1/(2n+1), then A+B =

(*)
If A,B are acute angles, tan A = n/(n+1), tan B = 1/(2n+1), then A+B = 

Grade:11

1 Answers

Yash Baheti IIT Roorkee
9 years ago
Hi,

Follow the steps to reach to the answer :

Apply the TAN (A+B) formula in terms of TAN (A) and TAN (B) and substitute the values. Upon simplification ull get TAN (A+B) = 1, i.e. A+B = pi/4.

This also goes with the fact that A and B are acute angles.